OpenStudy (alesia/bailey):

perimeter: find the value of x so that the triangles have the same perimeter p=5,x-5,x-4 p=16,10,x-10

OpenStudy (tylerd):

set them equal to eachother

OpenStudy (tylerd):

5+x-5+x-4 = 16+10+x-10

OpenStudy (tylerd):

2x-4=x+16

OpenStudy (alesia/bailey):

okay, we got that far, we got x=20 but they are not equal

OpenStudy (tylerd):

ok one sec

OpenStudy (aum):

Yes they are equal. With x = 20, the triangles are: 5, 15, 16 and 16, 10, 10. Both add up to 36 (perimeter)

OpenStudy (tylerd):

ya they are equal

OpenStudy (alesia/bailey):

omg nice thx

OpenStudy (aum):

The triangles are not equal but their perimeters are equal.

OpenStudy (aum):

yw.

OpenStudy (alesia/bailey):

ok so dnt have to plug the 20 bk n2 the equation?

OpenStudy (aum):

They are just asking for the x value and so you can stop it at x = 20. But for verification purposes you can put x back, find the perimeter of each triangle and make sure they are equal so you will know you have the right answer x = 20.